The first fellow could finish 7 jobs in 3 hours, and the second fellow could finish 8 jobs in 5 hours. How long would they have to work together to finish 59 jobs?
Answer by nerdybill(6962)

The first fellow could finish 7 jobs in 3 hours, and the second fellow could finish 8 jobs in 5 hours. How long would they have to work together to finish 59 jobs?
.
Let x = time (hours) to finish 59 jobs
then
x(7/3 + 8/5) = 59
multiplying both sides by 15:
x(35 + 24) = 885
x(59) = 885
x = 885/59
x = 15 hours